Home‎ > ‎Crew 133‎ > ‎


XO Gartrelle to CapCom:


The Greenhab is operating normally.

The fan is operating normally set to 68 degrees

The heater was turned down to 40 degrees at 9AM
The heater was turned up to 75 degrees at 5PM

The hydroponic vegetable seedlings were placed inside the Greenhab at 9AM
No growth was observed from the seedlings on this day. A small amount of water was added to the bottom of the seedling tray at about 10AM.
The seedlings were returned to the upper hab at 5PM. They will remain here for the night. 

All house plants were watered at approximately 10:15 AM
The herb seedlings and the Martian growth shoot were not watered in the morning because the soil appeared and felt sufficiently damp. The soil was checked again this afternoon and still felt damp enough that no water was needed. The soil will be checked in the AM for moisture. 

The cactus in the zen room was not watered. 

At 10:30 a dab of WD-40 was added to the front slide bolt on the Greenhab door to make it easier to open. The bolt is working properly at this time. 

Thermostat Readings from 5PM Today

                                         Current              Min                        Max
Temperature                  72 Degrees       41 Degrees           82 Degrees
Humidity                             16%                 14%                        21%